# CLUE: A Clinical Language Understanding Evaluation for LLMs

### v3

@@ 6.1 Effect of biomedical training
However, the superior performance of Mistral-7B-Instruct-v0.2, which shares the same architecture as BioMistral-7B-DARE, implies that an enhanced general domain pre-training dataset might have a more profound impact than fine-tuning alone. Assessing this conclusively is challenging, given the lack of details regarding the pre-training data for Mistral-7B-Instruct-v0.2.
@@
All biomedical models that showed improvement are based on the relatively low-scoring Mistral-7B-Instruct-v0.1. This leads us to our final hypothesis: these improvements might address gaps that more recent general-domain models, such as Meta-Llama-3 and Mistral-7B-Instruct-v0.2, have already overcome.
@@
6.2 Concerns About Data Contamination
Recently, data contamination has emerged as a significant concern in LLMs, particularly for open benchmarks like Huggingface’s Open LLM Leaderboard. This problem arises when the datasets used for evaluation have overlaps with training datasets, or when LLMs have been pre-trained or fine-tuned on the evaluation datasets themselves. Such contamination compromises the integrity of the evaluation process, skewing test results by favoring models for their memorization capabilities rather than their ability to generalize. Consequently, this has spurred a surge in research to address this issue (Li et al., 2024; Deng et al., 2023).
Accidental data contamination is less likely in datasets derived from sources such as MIMIC, where, for example, ICD-10 codes are not directly available but must be extracted separately. However, the risk of intentional benchmark manipulation persists. One approach to mitigate this is through the involvement of trusted entities, such as hospitals that have access to private data, to create their own confidential datasets. To foster the development of new datasets based on report summaries and allow users to assess LLMs with their private data, we will make our source code for the dataset generation process publicly available.
@@ 7 Conclusion
In conclusion, our study reveals that current biomedical LLMs are not competing effectively with general-domain models in clinical settings. While some biomedical models improved marginally, more recent models based on Meta-Llama-3 are underperforming. Fine-tuning these models with domain-specific data often results in decreased performance, introducing issues such as hallucinations and reduced model stability. This starkly contrasts traditional MCQA evaluations, such as USMLE, where biomedical models have previously shown superior performance. The introduction of the CLUE benchmark highlights these discrepancies, providing a more accurate and practical assessment of LLM capabilities in real-world healthcare applications. These findings underscore the need for ongoing evaluation and refinement to ensure the effective deployment of LLMs in medical practice.
@@ Appendix A Task Details
The characteristics of all included tasks are summarized in Table 1.
Dataset Samples Mean Words Metrics Focus
MedNLI 1425 20.81 Accuracy Clinical reasoning
MeQSum 1000 60.77 R-L, R-1, R-2, BERT F1 Summarization
LongHealth 400 5536.82 Accuracy Information extraction
Prob.Sum. 237 123.5 R-L, R-1, R-2, BERT F1, UMLS F1 Information extraction
Table 1: An overview of the characteristics of the tasks CLUE consists of.
@@ A.1 MeDiSumQA Generation
Mditron-7b llama2-7b Continued pretraining
@@
JSL-Med-Sft-Llama-3-8B Meta-Llama-3-8B-Instruct Unknown
JSL-MedLlama-3-8B-v1.0 Meta-Llama-3-8B-Instruct Unknown
JSL-MedLlama-3-8B-v2.0 Meta-Llama-3-8B-Instruct Unknown
@@ B.3 Prompting
Figures 9, 10, 11, 12, 13, and 14 are showing the prompt formats we are using for the different benchmark tasks. If the input length allowed this, we also included sample texts from the datasets.
@@ C.1 Error Analysis
llama2-7b Touvron et al. (2023a) 20.11 29.51 5.97 7.35 2.11 59.45 9.06 7.16 8.58 2.89 37.5
meditron-7b Chen et al. (2023) 11.46 2.39 6.49 7.87 2.59 59.57 12.52 6.17 7.47 2.53 40.53
llama2-7b-hf Touvron et al. (2023a) 36.94 41.27 17.43 22.25 6.93 66.33 21.62 36.45 39.99 18.13 75.98
Mistral-7B-v0.1 Jiang et al. (2023) 33.99 67.54 7.07 8.99 3.17 60.74 15.67 6.28 7.86 2.49 44.6
internistai/base-7b-v0.2 Griot et al. (2024) 25.98 50.00 4.67 6.08 1.59 57.87 6.8 5 5.85 0.74 38.54
BioMistral-7B Labrak et al. (2024) 40.79 62.32 18 21.85 8.63 66.96 20.92 24.58 27.38 12.1 67.05
BioMistral-7B-DARE Labrak et al. (2024) 43.24 66.76 18.81 22.92 8.93 68.93 22.65 26.16 29.69 13.49 68.68
Mistral-7B-Instruct-v0.1 Jiang et al. (2023) 40.31 64.79 14.46 18.6 6.26 65.79 20.08 21.85 25.26 11.17 66.15
zephyr-7b-beta Tunstall et al. (2023) 42.69 68.45 14.81 19.91 5.98 67.39 19.2 25.66 29.81 12.33 68.85
@@
Llama3-Aloe-8B-Alpha Gururajan et al. (2024) 42.50 73.94 9.47 12.3 4.06 65.56 15 23.78 27.02 12.31 66.01
JSL-Med-Sft-Llama-3-8B 19.00 29.08 4.59 5.75 1.61 57.69 8.75 6.57 7.73 2.99 31.70
JSL-MedLlama-3-8B-v1.0 17.00 13.31 12.87 15.40 4.73 65.18 18.42 6.02 7.19 2.43 41.78
JSL-MedLlama-3-8B-v2.0 31.26 35.70 8.16 10.61 2.94 64.48 13.17 30.93 34.54 16.31 71.03
Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024) 47.81 76.48 17.44 23.39 7.7 68.56 19.51 32.47 36.38 16.86 72.8
llama2-70b Touvron et al. (2023b) 35.17 76.27 7.13 8.77 3.15 59.61 14.34 3.75 4.15 1.05 33.59
meditron-70b Chen et al. (2023) 30.58 63.52 7.22 8.91 3.21 60.08 13.91 2.95 3.24 0.76 31.27
ClinicalCamel-70B Toma et al. (2023) 35.71 64.65 8.62 10.83 3.71 59.94 12.34 16.96 18.8 8.49 49.3
llama2-70b-hf Touvron et al. (2023b) 42.90 61.69 14.43 19.81 6.14 65.07 21.37 34.91 38.81 18.48 74.37
@@
Meta-Llama-3-8B-Instruct 40.61 68.3 66.55 56.25 22.44 28.15 9.63 68.62 22.74 3.95 17.55 61.93
@@
JSL-Med-Sft-Llama-3-8B 15.68 0.00 0.00 0.00 8.85 11.74 3.21 58.38 10.54 3.09 14.03 68.37
JSL-MedLlama-3-8B-v1.0 29.94 52.85 52.05 0.05 19.06 22.7 7.63 67.29 20.83 1.87 12.80 67.30
JSL-MedLlama-3-8B-v2.0 18.02 1.10 0.30 0.05 19.13 24.15 8.56 67.41 20.49 1.41 11.13 64.31

### v4

@@ 6 Discussion
This section explores the implications of our findings in comparison to existing evaluations, and the broader effects of biomedical training on model performance.
Model MCQA Level 1 Level 2
MEDITRON-7B +6.07 -7.08 -
MEDITRON-70B +3.63 -4.59 -
BioMistral-7B +4.13 +0.26 +0.71
BioMistral-7B-DARE +4.57 +2.93 +2.7
internistai/base-7b-v0.2 - -2.07 +5.52
Llama3-OpenBioLLM-8B -0.63 -15.17 -13.54
Llama3-OpenBioLLM-70B +1.46 -4.78 -10.45
Llama3-Med42-8B +0.47 +2.51 -1.4
Llama3-Med42-70B +2.8 -7.57 -15.14
Llama3-Aloe-8B-Alpha +2.21 -5.87 -8.67
Meditron3-8B - -2.76 –15.04
Meditron3-70B - -2.18 -8.51
Table 3: Reported performance difference after biomedical pretraining on MCQA tasks and our evaluation for models based on Llama-2-(7B/70B), Mistral-7B-v0.1, Meta-Llama-3-(8B/70B)-Instruct and Meta-Llama-3.1-(8B/70B)-Instruct.
6.1 Discrepancy to previous benchmarks
[…]
However, it is important to note that MCQA tasks primarily assess the medical knowledge of models without accounting for the challenges encountered in clinical practice, such as medical jargon, abbreviations, typos, irregular formatting, incomplete information, strict task formats, patient-friendly language, and long inputs. In contrast, our analysis, reveals a notable decline in performance after biomedical fine-tuning for many models. This decline suggests that our evaluation tasks, being more general and reflective of real-world scenarios, may expose deficiencies in the models’ retained general-domain capabilities following their specialization in biomedical tasks.
Notably, models based on the Mistral-7B architecture present mixed results, with some showing improvements while others experience performance deterioration. Despite these exceptions, the overall trend indicates a more pronounced decline in task performance than the improvements seen on MCQA benchmarks. This pattern suggests that biomedical fine-tuning may have an overall negative impact on the models’ ability to perform general clinical tasks, revealing a trade-off between domain-specific optimization and broader general-domain task performance.
@@ 6.2 Effect of biomedical training
However, the superior performance of Mistral-7B-Instruct-v0.2, which shares the same architecture as BioMistral-DARE, implies that an enhanced general domain pre-training dataset might have a more profound impact than fine-tuning alone.
@@
All biomedical models that showed improvement are based on the relatively low-scoring Mistral-7B-Instruct-v0.1. This leads us to our final hypothesis: these improvements might address gaps that more recent general-domain models, such as Llama-3 and Mistral-7B-Instruct-v0.2, have already overcome. This is supported by Tables 2 and 3 showing improvements of biomedical Mistral-7B models and performance deterioration of biomedical Llama3 models.
@@ 7 Conclusion
Our study reveals that biomedical LLMs are not competing effectively with general-domain models in practical medical settings. While some biomedical models have shown marginal improvements, many recent models are underperforming. Fine-tuning these models with domain-specific data often leads to reduced performance, introducing hallucinations and decreased model stability. This stands in contrast to traditional MCQA evaluations, where biomedical models have previously demonstrated superior performance. Our evaluation provides a more practical and accurate assessment of LLM capabilities in real-world healthcare settings. To support further progress in this field, we are open-sourcing our evaluation scripts, allowing for broader validation and replication of these results. We believe future research should focus on addressing the highlighted issues to enhance the effective deployment of LLMs in medical practice.
@@ A.1 MeDiSumQA Generation
Meditron-7B llama2-7b Continued pretraining
@@
Llama3-Med42-8B Meta-Llama-3-8B-Instruct SFT + DPO
@@
Llama3-Med42-70B Meta-Llama-3-8B-Instruct SFT + DPO
Meditron3-70B Meta-Llama-3.1-70B-Instruct -
@@ B.3 Prompting
Figures 8, 9, 10, 11, 12, and 13 are showing the prompt formats we are using for the different benchmark tasks. If the input length allowed this, we also included sample texts from the datasets.
@@ C.1 Error Analysis
Llama-2-7b 20.11 29.51 5.97 7.35 2.11 59.45 9.06 7.16 8.58 2.89 37.5
Meditron-7B 13.03 2.39 11.29 13.43 4.81 63.35 15.21 6.83 7.90 2.26 43.39
Llama-2-70b 35.17 76.27 7.13 8.77 3.15 59.61 14.34 3.75 4.15 1.05 33.59
Meditron-70B 30.58 63.52 7.22 8.91 3.21 60.08 13.91 2.95 3.24 0.76 31.27
Mistral-7B-v0.1 33.99 67.54 7.07 8.99 3.17 60.74 15.67 6.28 7.86 2.49 44.6
Mistral-7B-Instruct-v0.1 40.31 64.79 14.46 18.6 6.26 65.79 20.08 21.85 25.26 11.17 66.15
zephyr-7b-beta 42.69 68.45 14.81 19.91 5.98 67.39 19.2 25.66 29.81 12.33 68.85
BioMistral-7B 40.57 62.75 16.9 20.89 8.4 59.03 20.12 25.89 28.46 13.31 67.93
BioMistral-7B-DARE 43.24 66.76 18.81 22.92 8.93 68.93 22.65 26.16 29.69 13.49 68.68
internistai/base-7b-v0.2 38.24 76.34 13.11 16.79 5.63 62.94 17.22 7.65 9.54 3.44 40.34
@@
Llama3-Med42-8B 50.88 77.46 24.43 29.18 10.76 72.51 25.12 36.03 40.32 19.83 74.93
Llama3-Aloe-8B-Alpha 42.50 73.94 9.47 12.3 4.06 65.56 15 23.78 27.02 12.31 66.01
Meta-Llama-3.1-8B-Instruct 50.32 79.08 20 26.14 8.53 70.79 23.5 35.69 39.55 18.11 75.03
Meditron3-8B 47.56 74.01 18 22.39 8.1 69.26 21.8 33.61 37.6 17.84 74.01
@@
Meta-Llama-3-8B-Instruct 38.98 68.3 66.55 41.6 22.44 28.15 9.63 68.62 22.74 3.95 17.55 61.93
@@
Meta-Llama-3.1-8B-Instruct 45.61 76.65 75.4 59.45 26.23 31.52 11.45 70.38 24.82 6.72 22.54 71.08
Meditron3-8B 30.57 73.65 70.5 7.3 23.52 28.71 10.61 69.1 23.63 2.23 6.25 21.89